Solve the equation and find x for the following equation:- 6-x/x2-4=2+ x/x+2

Grade:8

1 Answers

Arun
5 years ago
Dear Aniruddh
Please take care of brackets
anyways I am solving it what I am thinking of – 
(6 – x) / (x^2 – 4) = 2 + x / (x + 2)
(6 – x) / (x – 2) = 2x + 4 + x
6 – x = (x – 2) (3x +4)
6 – x = 3x^2 – 2x – 8
3x^2 – x – 14 = 0
3x^2  + 6x – 7x – 14 = 0
3x (x + 2) – 7 (x + 2) = 0
x = – 2 and x = 7/3
 but x = – 2 is not acceptable 
hence x = 7/3 is answer